The area of a triangle formed by the intersection of a line parallel to x-axis and passing through P(h,k) with lines y=x and x+y=2 is 4h^2 then locus of P is ??

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Area of triangle = 1/2(AB. AC)=4h2
and AB = \sqrt{2}|k – 1| = AC
⇒ 4h2=(1/2)(2)(k-1)2
⇒ k – 1 = ± 2h.
⇒ locus is y = 2x + 1, y = – 2x + 1.
Thanks and Regards
Shaik Aasif
askIITians faculty